OpenStudy (anonymous):

completing the square x^2-4x+1=0

OpenStudy (lgbasallote):

x^2 - 4x + 4 = -1 +4 (x-2)^2 = 3 x-2 = sqrt of 3 x = (sqrt of 3) + 2

OpenStudy (anonymous):

no, the ans should be (x - 2)^2 + 3

hero (hero):

x^2 -4x+ 1 = 0 x^2 -4x = -1 x^2 - 4x + 4 = -1 +4 (x-2)^2 = 3 x-2 = +/-sqrt{3} x = +/-sqrt{3} + 2

hero (hero):

tanvidais 13, you have to solve for x when you complete the square